What is the probability of rolling and even number on an ordinary dice?
your set of even number x = {2,4,6}, which tells you that 3 out the 6 sides are even. Your probability is 3/6 = 1/2
What are the 4 r's to prevent head-on collison?
How about r for reverse. If you are always moving in reverse you can never have a head on collision. Simple!
4/7 * 4/7 or 16/49.
What is the probability that the upturned face is 3 or less?
The answer will depend on what the experiment is.
Let the number of sweets be n and use the rules of probability:-
If: 6/n times 5/n-1 = 1/8
Then: 30/n2-n = 1/8
Multiplying both sides by n2-n: 30 = n2-n/8
Multiplying both sides by 8: 240 = n2-n
Subtracting both sides by 240: 0 = n2-n-240
Solving the above quadratic equation: n = 16 or n = -15
Therefore it follows: n = 16 sweets
Note that n2 means n squared
What is the exponent of ten beyond which statistical probability becomes zero?
Although it's not clear which distribution you're referring to, I sense that I know what you mean. This is one of those situations where a function is only asymptotically zero, which is to say that it reaches zero only in the limit. But, no matter how large the exponent is, it never actually reaches zero; therefore, the probability never becomes zero. It only becomes extremely small.
What is the probability of rolling two dice and getting the same number?
1/6 is the answer
Another Answer:-
It is: 1/6*1/6 = 1/36
Will running more trials affect results of a sampling distribution?
Yes. It should reduce the variance.
It is 15/50 = 0.3
If they can all be distiguished from one another, then 8*2 = 16 ways.
What is the probability associated with the area to the right of the mean in a normal distribution?
50%
What is the probability of getting a sum of 2?
That would depend on the conditions of the experiment.
There are (1*5*4)*(3*2*1) = 120 arrangements.
55mph
Whats is a good black and red deck?
A good red black deck has a lot of creatures that deal lots of damage. The black portion of the deck should have some zombies and other undead stuff. The red portion should be able to boost power and deal a lot of damage
The probability is 1. If the maximum number of hairs is 100000 then there can be only 100000 people, each with a different number of hairs: 1, 2, 3, ..., 100000. Then the 100001st person must have the same number as someone in that list. Since the population of NYC is greater than 100001, then the event is a certainty.